OpenStudy (anonymous):

Help with geometric sequences and series please!! Need to know how to solve this problem!

OpenStudy (damonte212):

whats the problem

OpenStudy (anonymous):

sorry it wasnt posting before!

OpenStudy (anonymous):

I am really lost on how to solve it. do you know how to do questions like this?

OpenStudy (anonymous):

@hartnn are you able to help with this?

OpenStudy (anonymous):

@Abhisar any help?

OpenStudy (anonymous):

@dan815 sorry for just tagging random people but i cant figure it out!

OpenStudy (dan815):

ok

OpenStudy (dan815):

u can factor out the 4

OpenStudy (dan815):

sum 4* 2^i=4*sum 2^i sum 2^i = 2+4+8....+2^9+2^10 2*sum 2^i-sum2^i = 2^11-2 sum2^i(2-1)=2^11-2 sum2^i=2^11-2 4*sum2^i=4*(2^11-2)

OpenStudy (anonymous):

im not sure thats what we are learning hahaha that looks very different

OpenStudy (dan815):

i can go step by step

OpenStudy (dan815):

if u want to learn

OpenStudy (anonymous):

we usually put it into a formula like the sum of infinite series

OpenStudy (anonymous):

or sum of finite series

OpenStudy (anonymous):

like these.

OpenStudy (dan815):

let S be that expression they gave u S=4*2^1+4*2^2+4*2^3....+4*2^10

OpenStudy (dan815):

does this make sense so far

OpenStudy (anonymous):

oops i attached the wrong picture. but yes somewhat

OpenStudy (dan815):

i am rewriting the summation notation in terms

OpenStudy (dan815):

and since 4=2^2

OpenStudy (dan815):

we can rewrite it as S=4*2^1+4*2^2+4*2^3....+4*2^10 =2^3+2^4+2^5......+2^12

OpenStudy (dan815):

make sense so far?

OpenStudy (anonymous):

yes i think im following

OpenStudy (anonymous):

sorry my computer has issues with openstudy! got booted off for a second.

OpenStudy (anonymous):

here is the picture i was trying to attach before.

OpenStudy (dan815):

okay now we can do a trick here

OpenStudy (dan815):

S=2^3+2^4+2^5......+2^12 2*S=2^4+2^5+2^6....+2^13 2*S-S=2^13-2^3 S(2-1)=2^13-2^3 S=2^13-2^3

OpenStudy (anonymous):

okay i believe i get what you are saying. but that cant be my answer can it?

OpenStudy (dan815):

S=a^1+a^2+a^3......+a^k a*S=a^2+a^3+a^4....+2^(k+1) a*S-S=a^(k+1)-a S(a-1)=a^(k+1)-a S=(a^(k+1)-a)/(a-1)

OpenStudy (dan815):

that is the formula u put up there

OpenStudy (dan815):

it comes from the same method

OpenStudy (dan815):

in our case 2 is the geometric factor if the factor is changed to r then u get that same formula

OpenStudy (dan815):

just see if u can figure this out, u can forget the geometric series formula if u understand this http://prntscr.com/5pno2t

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

@cwrw238 any ideas?

OpenStudy (anonymous):

@SolomonZelman